4. Contribution Margin Ratio

a. Young Company budgets sales of $1,080,000, fixed costs of $43,700, and variable costs of $194,400. What is the contribution margin ratio for Young Company?

_______%

b. If the contribution margin ratio for Martinez Company is 63%, sales were $556,000, and fixed costs were $269,720, what was the operating income?

$

5. Break-even sales and sales to realize operating income

For the current year ended March 31, Cosgrove Company expects fixed costs of $494,400, a unit variable cost of $49, and a unit selling price of $73.

a. Compute the anticipated break-even sales (units).
units

b. Compute the sales (units) required to realize operating income of $112,800.
units
Answer #1

4) Contribution margin ratio = (1080000-194400)/1080000 = 82%

b) Operating income = (556000*63%)-269720 = 80560

5a) Break even unit = 494400/(73-49) = 20600 Units

5b) Required unit = (494400+112800)/24 = 25300 Units